OpenStudy (anonymous):

Four friends are collecting canned goods for charity. Ron collects 10 more than 3 times as many cans as Jasmine. Barry collects twice as many cans as Jasmine. Alicia collects 10 more cans than Barry. How many cans does each person collect if Jasmine and Ron collect the same number of cans as Alicia and Barry?

OpenStudy (anonymous):

At least to clear it up: ron = r jasmine = j barry = b alicia = a Ron collects 10 more than 3 times as many cans as Jasmine... r = 10 + (3j) Barry collects twice as many cans as Jasmine... b = 2j Alicia collects 10 more cans than Barry... a = 10 + b Jasmine and Ron collect the same number of cans as Alicia and Barry... jr = ab I'm not sure if you could use substitution here.

hartnn (hartnn):

'Jasmine and Ron collect the same number of cans as Alicia and Barry... jr = ab'<---no. Jasmine and Ron collect the same number of cans as Alicia and Barry... j+r = a+b

OpenStudy (anonymous):

^Exactly, I made a mistake there.... it's j + r = a + b